OpenStudy (samigupta8):

the interval for which the series 1+(x-1)+(x-1)^2+.........+≠ may be summed is

OpenStudy (samigupta8):

it's upto infinity

OpenStudy (amistre64):

\[\sum(x-1)^n\] \[\lim_{n\to inf}\frac{(x-1)^{n+1}}{(x-1)^n}\] \[\lim_{n\to inf}(x-1)\] \[|x-1|\lim_{n\to inf}1=|x-1|\] \[\|x-1|<1\] \[-1<x-1<1\] \[0<x<2\]